if: C = 120 + 0.8Y, I = 200 + 0.05i and Y = 1400, What is the size of government expenditure multiplier and savings?
1
Expert's answer
2015-03-19T11:32:25-0400
We can use the following formulas:
Government expenditure multiplier = 1 / (1-MPC)
MPC is the marginal propensity to consume
Government expenditure multiplier = 1 / (1-0.8) = 5
We have the consumption function in the form C = a + MPC (Y) that
autonomous consumption is a. The savings function is S = -a + (1 –
MPS) (Y)
So, the saving function is:
S = -120 + 0.2 Y
S = -120 + 280 = 160